equation of the diameter of the circle is given by x2 + y2 -12x -4y +6 =0 is given by?

Grade:12th pass

1 Answers

Waseem khan
6 years ago
FROM HERE WE GET THAT THE CENTRE IS (6,-4) SO RHE DIAMETER WILL PASS THROUGH (6,-4) AND HERE THE SLOPE IS NOT GIVEN TO US AND NOT A POINT SO WE CAN TAKE ANY SLOPE LET US TAKE A SLOPE 1 (Y+4)=1(X-6)X-Y=10SO WE CAN GET INFINITELY MANY EQUATIONS